OpenStudy (anonymous):

solve for y: 8y^2=x(x-4)^2

OpenStudy (anonymous):

okay

OpenStudy (confluxepic):

First distribute.

OpenStudy (anonymous):

my final answer was y=x-4 square root of 8x over 8

OpenStudy (confluxepic):

Then you divide by 8 on both sides.

OpenStudy (anonymous):

but im not sure

OpenStudy (anonymous):

you got that?

OpenStudy (campbell_st):

there is no need to distribute divide both sides by 8 then take the square root \[y = \sqrt{\frac{x(x-4)^2}{8}}\]

OpenStudy (campbell_st):

you can simplify it from here if needed.

OpenStudy (anonymous):

and when you simplify

OpenStudy (anonymous):

you have y=x-4 square root of 8x over 8?